OpenStudy (anonymous):

Solve using elimination: -9x = -4y − 10 -11x + 5y = -13

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

First, write both equations in the form \(ax + by = c\) The second is already in that form.

OpenStudy (anonymous):

so it -9x+-4y=-10

OpenStudy (mathstudent55):

No. You need to add 4y to both sides: -9x = -4y − 10 +4y +4y --------------------add -9x + 4y = -10 Now write one above the other.

OpenStudy (anonymous):

ok i did that now what do i do

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

Now we have this. -9x + 4y = -10 -11x + 5y = -13

OpenStudy (anonymous):

yes and i got this -20x+9y=-23

OpenStudy (mathstudent55):

Now we need to find which variable to eliminate. Since by simply adding the equations neither variable gets eliminated, we need to do what we did last time. We have to multiply both equations by numbers that will make one variable add to zero.

OpenStudy (mathstudent55):

As you can see, neither variable was eliminated when you added the equations.

OpenStudy (anonymous):

ok so what do i multiply the first equestion by

OpenStudy (mathstudent55):

-9x + 4y = -10 -11x + 5y = -13 Look at the y terms. If you multiply the first equation by -5, you get -20y. If you multiply the second equation by 4, you get 20y. -20y and 20y add up to zero. That would eliminate the y variable, just what we need.

OpenStudy (anonymous):

then do i add them to elimanate them

OpenStudy (mathstudent55):

Yes, but first multiply them by the numbers.

OpenStudy (anonymous):

x=-2

OpenStudy (anonymous):

the answers are x=-2 y=-7

OpenStudy (mathstudent55):

45x - 20y = 50 -44x + 20y = -52 --------------------add x = -2 You are correct. Now insert -2 for x in one of the original equations, and solve for y.

OpenStudy (anonymous):

i got it ima try the next one on my own and if i need your help i'll ask you if that ok

OpenStudy (mathstudent55):

What did you get for y?

OpenStudy (mathstudent55):

-9x = -4y − 10 -9(-2) = -4y - 10 18 = -4y - 10 28 = -4y y = -7

OpenStudy (anonymous):

i got -7

OpenStudy (mathstudent55):

Great.

OpenStudy (anonymous):

can you help me on one i just posted